# Characterizing NGC 6383: A study of pre-main sequence stars, mass segregation, and age using Gaia DR3 and 2MASS

### v1

@@ 3.5 Comparative analysis with cluster members and PMS study
3.5.1 Peculiar Stars
NGC 6383 22  [^3] Cataloged in Simbad as Gaia DR3 4054615634716139264, is identified as a $\lambda$ Bootis star. These stars are chemically peculiar A-type stars with abundance anomalies attributed to the accretion of metal-poor material. In a spectroscopic survey of metal-weak and emission-line stars in the Southern Hemisphere, Murphy et al. (2020) found that $\lambda$ Boo stars, including NGC 6383 22, exhibit excesses at longer wavelengths, suggestive of circumstellar discs. This characteristic aligns with other stars in NGC 6383 and those studied in Kalari (2019). Unfortunately, NGC 6383 22, despite matching the position and parallax of the cluster, does not belong to NGC 6383 due to its differing proper motion, indicating it is passing near or through the cluster.
Figure 15: Radial velocity plotted against G-band magnitude for NGC 6383 sources categorized by the presence of robust radial velocity amplitude data from Gaia. This measure reflects the total amplitude in the radial velocity time series following outlier removal. Yellow circles indicate sources without amplitude measurements, displaying radial velocity uncertainties with error bars. Blue circles represent sources with amplitude data, color-coded by amplitude value according to the scale on the right. This representation allows for detailed analysis of amplitude variations relative to radial velocities and magnitudes in the cluster.
@@ 3.5.2 HD 159176
HD 159176, a double-lined spectroscopic binary composed of O7V type stars, is located in the projected center of NGC 6383 and is responsible for ionizing the surrounding HII region (Penny et al., 2016). The estimated age of HD 159176 is around $2.3-2.8$ Myr (Rauw et al., 2010). Previous analyses suggested that HD 159176 might predate the cluster itself and could have initiated star formation in the cluster’s core and surrounding areas (Fitzgerald et al., 1978).
In the study by Aidelman et al. (2018), it was hypothesized that if HD 159176 were a blue straggler, it would imply an older age for NGC 6383, ranging between $6$ and $10$ Myr. However, our findings indicate that HD 159176 is not a member of NGC 6383. The proper motion of HD 159176 ($\mu_{\alpha}^{*}=2.621\pm 0.010$ mas yr-1, $\mu_{\delta}=-0.797\pm 0.009$ mas yr-1) differs from the mean proper motion of the cluster. Additionally, the parallax of HD 159176 ($1.153\pm 0.024$ mas) is outside the $2\sigma$ range of the cluster’s parallax distribution.
Despite its central position, these discrepancies in proper motion and parallax confirm that HD 159176 is not gravitationally bound to NGC 6383. Therefore, the hypothesis that HD 159176 is a blue straggler influencing the cluster’s age must be reconsidered. Our study supports the notion that NGC 6383 has an age of approximately 4 Myr, independent of HD 159176’s characteristics.
@@ 3.5.3 PMS Study Using VPHAS+
Kalari (2019) examined 55 classical T-Tauri stars (CTTS) in the star-forming region Sh 2-012 and NGC 6383, utilizing optical photometry and Gaia astrometry. They reported a median age of $2.8\pm 1.6\,\mathrm{Myr}$ for these stars, with masses ranging from $0.3$ to $1\,M_{\odot}$. Notably, $94\%$ of CTTS with near-infrared cross-matches conform to the near-infrared T-Tauri locus, and all stars with mid-infrared photometry showed signs of accreting circumstellar discs concentrated around NGC 6383.
Out of the 55 CTTS, only 15 are catalogued as members or potential members in this study, with most having a PMS probability over $0.6$  [^4] Except for Gaia DR3 4054615467234669440, Gaia DR3 4054617116502154368, Gaia DR3 4054618761456141056, and Gaia DR3 4054826809671572096, which have lower probabilities of $16$, $12$, $55$, and $45$ per cent respectively. Only one T-Tauri source  [^5] Catalogued as Gaia DR3 4054567805963940352 is likely a binary, with a $0.82$ probability of being a binary star according to our findings.
3.5.4 Catalogs Comparison
We compared our membership list with the catalogs from Cantat-Gaudin et al. (2020), Jaehnig et al. (2021), He et al. (2022), Hunt & Reffert (2024), and the members listed in SIMBAD.
From Cantat-Gaudin et al. (2020), we identified 148 common sources, found 97 sources not cross-matched in their catalog, and 173 sources unique to our data. Jaehnig et al. (2021) showed 161 sources in common, with 123 sources not present in our data, and 160 not listed in their catalog. The catalog of He et al. (2022) includes 90 common sources, 47 not found in our data, and 231 not in their catalog. Hunt & Reffert (2024) had 202 common sources, with 120 unique to their catalog, and 119 unique to our results. Using the SIMBAD database, we confirmed 168 common sources, with 163 unique to SIMBAD and 153 unique to our results.
@@ 4 Conclusion
This study provides a comprehensive analysis of the open cluster NGC 6383, leveraging a combination of Bayesian analysis and machine learning techniques to identify cluster members and determine key parameters. Using the HDBSCAN algorithm, we identified 321 likely cluster members with membership probabilities exceeding 60 percent.
@@
Our findings indicate a mean cluster age of approximately $4~{}\mathrm{Myr}$, with a distance of $1.207~{}\mathrm{kpc}$ derived from both parallax and distance modulus methods. We determined the cluster’s core and tidal radii to be $1.95~{}\mathrm{arcmin}$ and $40.53\mathrm{arcmin}$, respectively, and observed significant mass segregation, particularly among binary stars.
@@
The CMD analysis revealed a well-defined main sequence and a population of PMS stars, highlighting the cluster’s recent star formation activity. The Sagitta neural network provided further insights into PMS probabilities, ages, and extinctions of the cluster members.
Comparison with historical data and previous studies shows our results align well with recent estimates, offering updated parameters for NGC 6383. The cluster’s structural and dynamic properties, including proper motions, radial velocities, and spatial distribution, were thoroughly characterized, contributing valuable data for future research on young open clusters.
@@
We observed that the binary stars are more centrally concentrated compared to single stars, suggesting both dynamical and primordial mass segregation. The calculated half-mass relaxation time is approximately $9.8~{}\mathrm{Myr}$, indicating that NGC 6383 is not yet fully relaxed. The observed central concentration of binaries, even for those with segregation times longer than the cluster’s age, suggests that these stars may have formed closer to the cluster center or within rapidly evolving clumps that merged early in the cluster’s history.
@@
Overall, this study demonstrates the effectiveness of combining advanced computational techniques with traditional astronomical methods to enhance our understanding of stellar clusters. The detailed characterization of NGC 6383 provides a solid foundation for further investigations into its formation history and evolution.
Acknowledgements
We gratefully acknowledge support by the ANID BASAL project FB210003 and the SOCHIAS GEMINI project 32230014.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ Acknowledgements
This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work made use of Astropy: [^6] http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022).
Data Availability
Appendix A Some extra material
@@
Figure 16: Cluster size variation as a function of minimum cluster size, determined through HDBSCAN clustering on proper motion data of sources. The blue line represents cluster sizes achieved at various minimum cluster sizes. The dashed blue line highlights the optimal minimum cluster size of 43, where the cluster size reaches a peak. The dashed green line indicates the maximum observed cluster size of 701.

### v2

@@ 3.5.1 HD 159176
HD 159176, a double-lined spectroscopic binary composed of O7V type stars, is located in the projected center of NGC 6383 and is responsible for ionizing the surrounding HII region (Penny et al., 2016). The estimated age of HD 159176 is $2.30-2.80$ Myr (Rauw et al., 2010). Previous analyses suggested that HD 159176 might predate the cluster itself and could have initiated star formation in the cluster’s core and surrounding areas (Fitzgerald et al., 1978).
In the study by Aidelman et al. (2018), it was hypothesized that if HD 159176 were a blue straggler, it would imply an older age for NGC 6383, ranging between $6.00$ and $10.0$ Myr. However, our analysis indicates that HD 159176 is not a member of NGC 6383. The proper motion ($\mu_{\alpha}^{*}=2.620\pm 0.010$ mas yr-1, $\mu_{\delta}=-0.7970\pm 0.0090$ mas yr-1) and parallax ($1.150\pm 0.024$ mas) are outside the $3\sigma$ width of the cluster posterior distribution of $\mu_{\delta}$ and $\varpi$, whose distributions can be seen in Fig. 3 and Fig. 4.
Despite its central position, these discrepancies in proper motion and parallax confirm that HD 159176 is not gravitationally bound to NGC 6383. Therefore, the hypothesis that HD 159176 is a blue straggler influencing the cluster’s age must be reconsidered. Our study supports the notion that NGC 6383 has an age of approximately 4 Myr, independent of the characteristics of HD159176.
3.5.2 NGC 6383 22
NGC 6383 22333Cataloged in Simbad as Gaia DR3 4054615634716139264, is identified as a $\lambda$ Bootis star. These stars are chemically peculiar A-type stars with abundance anomalies attributed to the accretion of metal-poor material. In a spectroscopic survey of metal-weak and emission-line stars in the Southern Hemisphere, Murphy et al. (2020) found that $\lambda$ Boo stars, including NGC 6383 22, exhibit excesses at longer wavelengths, suggestive of circumstellar discs. This characteristic aligns with other stars in NGC 6383 and those studied in Kalari (2019). However, NGC 6383 22, despite matching the position and parallax of the cluster, does not belong to it because of its proper motion ($\mu_{\alpha}^{*}=1.790\pm 0.030$ mas yr-1, $\mu_{\delta}=-2.710\pm 0.024$ mas yr-1), which is more than $3\sigma$ away from the mean of the posterior distribution of $\mu$. This result indicates that this star is likely passing near or through the cluster.
@@ 3.5.3 PMS Study Using VPHAS+
Kalari (2019) examined 55 classical T-Tauri stars (CTTS) in the star-forming region Sh 2-012 and NGC 6383, utilizing optical photometry and Gaia astrometry. They reported a median age of $2.80\pm 1.60\,\mathrm{Myr}$ for these stars, with masses ranging from $0.3$ to $1.0\,M_{\odot}$. Notably, $94\%$ of CTTS with near-infrared cross-matches conform to the near-infrared T-Tauri locus, and all stars with mid-infrared photometry showed signs of accreting circumstellar discs concentrated around NGC 6383.
Out of the 55 CTTS kinematic members of Sh 2-012, only 15 are catalogued as members or potential members of the cluster in this study, with most having a PMS probability over 0.60.60.60.6444Except for Gaia DR3 4054615467234669440, Gaia DR3 4054617116502154368, Gaia DR3 4054618761456141056, and Gaia DR3 4054826809671572096, which have probabilities of 16, 12, 55, and 45 percent, respectively.. The difference in member selection between the studies is primarily due to the differences in the membership determination methods. It should be noted that Kalari (2019) determined the membership of NGC 6383 by modeling the distribution of proper motions with a double-peaked Gaussian. Thus, the difference between this approach and the one adopted in the present paper surely induced a difference between the samples of cluster members selected in the two works. Only one T-Tauri source [^5] Catalogued as Gaia DR3 4054567805963940352 is likely a binary, with a $0.82$ probability of being a binary star according to our analysis.
3.5.4 Comparison with published catalogs
We compared our membership list with the catalogs from Cantat-Gaudin et al. (2020), Jaehnig et al. (2021), He et al. (2022), Hunt & Reffert (2024), and the members listed in SIMBAD. From Cantat-Gaudin et al. (2020), we identified 148 common sources, found 97 sources not cross-matched in their catalog, and 173 sources unique to our selection. On the other hand, we find 161 matches with the sample of Jaehnig et al. (2021), with 123 sources not present in our data, and 160 not listed in their catalog. The catalog of He et al. (2022) includes 90 common sources, 47 not found in our data, and 231 not in their catalog. Hunt & Reffert (2024) had 202 common sources, with 120 unique to their catalog, and 119 unique to our results. Using the SIMBAD database, we confirmed 168 common sources, with 163 unique to SIMBAD and 153 unique to our sample. It should be noted that all of these studies used Gaia data.
@@ 4 Conclusion
This study presents an analysis of the young open cluster NGC 6383, employing Bayesian analysis and machine learning techniques to identify cluster members and determine key parameters. Using the HDBSCAN algorithm, we identified 254 probable cluster members with membership probabilities exceeding 60 percent.
@@
Our findings indicate that the cluster has a mode age of $3.53^{+1.40}_{-1.00}\leavevmode\nobreak\ \mathrm{Myr}$, with a distance of $1.110\pm 0.060\,\mathrm{kpc}$ derived from the parallax-based distance estimation. The core and tidal radii of the cluster were determined to be $1.950\pm 0.190\,\mathrm{arcmin}$ and $40.4\pm 14.3\,\mathrm{arcmin}$, respectively. We also observed mass segregation among binary stars, indicating that NGC 6383 is not fully relaxed.
@@
The CMD analysis revealed a well-defined main sequence and a population of PMS stars, indicating an age spread of star formation within the cluster ranging from $1$ to $6$ Myr ago, suggesting ongoing star formation. The Sagitta neural network provided additional insights into the probabilities of PMS stars, their ages, and the extinction affecting cluster members.
Our analysis aligns with recent studies, offering updated parameters for NGC 6383. A significant result is the identification of HD 159176 as a non-member of NGC 6383. Despite its central position and previous assumptions about its influence on the cluster’s formation, our analysis of proper motion and parallax data indicates that HD 159176 is not gravitationally bound to NGC 6383. This has important implications for determining the cluster’s age, now estimated independently of this source’s characteristics.
@@
Additionally, we observed that binary stars are more centrally concentrated compared to single stars, suggesting primordial mass segregation. The calculated half-mass relaxation time indicates that NGC 6383 is not yet fully relaxed. The observed central concentration of binaries, even for those with segregation times longer than the cluster’s age, suggests these stars may have formed closer to the cluster center. This finding is consistent with other studies and simulations where initial mass segregation and gas expulsion play critical roles in star cluster evolution, which is in agreement with the studies of McMillan et al. (2007); Marks et al. (2008); Chen et al. (2007); Dib et al. (2007); Sabbi et al. (2008). Regardless of the specific mechanism, this suggests a primordial, or near-primordial, central concentration of binaries during the formation process of the cluster, which is consistent with other young open clusters in those studies and simulations.
@@
Overall, this study demonstrates the effectiveness of combining advanced computational techniques with traditional astronomical methods to enhance our understanding of stellar clusters. The characterization of NGC 6383 provides a solid foundation for future investigations into its formation history and evolution. This study is a precursor for future work investigating mass segregation, pre-main sequence stars, age, and membership of new open cluster candidates in the Gaia era. The software used in this study is available on GitHub [^6] https://github.com/notluquis/COSMIC, and the catalog will be archived at the CDS.
Acknowledgements.
@@
We gratefully acknowledge support by the ANID BASAL project FB210003 and the SOCHIAS GEMINI project 32230014. L.P. gratefully acknowledges the invaluable comments and guidance of Dr. Robert Mathieu in the preparation of this investigation. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work made use of Astropy: [^7] http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022).
Appendix A HDBSCAN diagnostic and ASteCA posterior corner plots
The present section describes the runs of HDBSCAN and presents the resulting plots, providing a comprehensive visualization of the clustering process. These figures illustrate the cluster size variation as a function of minimum cluster size, the hierarchical structure of the cluster system through a condensed cluster tree, and the posterior distributions of inferred parameters.
@@ Appendix A HDBSCAN diagnostic and ASteCA posterior corner plots
Figure 16: Cluster size variation as a function of minimum cluster size, determined through HDBSCAN clustering on the proper motion data of the sources. The blue line represents cluster sizes achieved at various minimum cluster sizes. The dashed blue line highlights the optimal minimum cluster size of 43, where the cluster size reaches a peak. The dashed green line indicates the maximum observed cluster size of 701.
Figure 16 shows how the size of clusters varies with different minimum cluster size parameters in HDBSCAN. This plot helps visualizing the optimal minimum cluster size, marked by the dashed blue line, where the cluster size reaches its maximum efficiency. The dashed green line represents the maximum observed cluster size, providing a benchmark for comparison.
@@
Figure 17 presents the condensed cluster tree, which visually represents the hierarchical structure of the clusters. This dendrogram highlights the proximity or potential mergers of clusters and their degree of separation. The left side of the plot focuses on the NGC 6383 cluster sources, while the right side shows the HDBSCAN-identified field sources. The color bar indicates the number of sources at each hierarchical level, with $\lambda$ values representing the inverse of the distance.
@@
Figure 18 displays the posterior distributions of key parameters: visual extinction ($A_{v}$), distance modulus ($dm$), logarithmic age ($loga$), and metallicity ($met$). The diagonal plots show the marginal distributions for each parameter, providing insights into their individual probability densities. The off-diagonal hexbin plots illustrate the joint distributions, in which it is possible to see correlations between the parameters. The black lines indicate the mode of the distributions, highlighting the most probable values.